OpenStudy (anonymous):

---NEED HELP!--- Henry can paddle his kayak in still water with a speed of 20m per minute. If Henry paddles 100m against the current in less time as he paddles 150m with the current. Find the speed of the current. Assume that at no time is he being swept away by the current.

OpenStudy (anonymous):

@satellite73 @.Sam. @Callisto @amistre64 @myininaya @robtobey

OpenStudy (anonymous):

Help please?

OpenStudy (zpupster):

20+c downstream 20-c upstream c=current 20m/min=rate of the kayak in still water then 100(20+c)=150(20-c ) 2000+100C=3000-150C

OpenStudy (anonymous):

C=4

OpenStudy (anonymous):

This is an inequality question @zpupster @MajicMuzyk.

OpenStudy (anonymous):

@zpupster I don't get this part 20+c downstream 20-c upstream c=current

OpenStudy (anonymous):

and the part after @zpupster

OpenStudy (anonymous):

why did you multiply distance by speed?

OpenStudy (anonymous):

Can you also answer this question: factorise 3x^2 - 40 + 50.

OpenStudy (zpupster):

3x2+10 that is factored for the other 20 + c- what the current is downstream + what he paddles in still H2O and 20 -c the current upstream is neg,

OpenStudy (zpupster):

part after goes like this 2000+100C=3000-150C add 150c to both sides and subtract 2000 250c = 1000 divide both sides by 250 c=4 m/min

OpenStudy (anonymous):

I forgot the x sorry 3x^2-40x+50

OpenStudy (anonymous):

but the answer is o<c<4???

OpenStudy (anonymous):

m/min

OpenStudy (anonymous):

Why did you multiply here? 100(20+c)=150(20-c )